# SPEAR: A Unified SSL Framework for Learning Speech and Audio Representations

### v1

@@ Appendix F HEAR Evaluation
[…]
| SPEAR, Base | 95M | 92.0 | 78.6 | 93.8 | 83.8 | 49.9 | 86.5 | 96.9 | 95.2 | 70.8 | 24.7 | 96.8 | 94.1 | 78.9 | 64.6 | 97.0 | 96.7 | 21.9 | 77.3 | 77.8 |
@@
| SPEAR, XLarge | 600M | 94.5 | 81.6 | 95.5 | 84.8 | 52.4 | 88.5 | 98.5 | 94.4 | 77.7 | 27.7 | 97.0 | 96.5 | 81.5 | 63.4 | 97.2 | 98.1 | 22.6 | 83.6 | 79.7 |
[…]
@@
By performing unified speech and audio pre-training that incorporates more speech data, the dual-domain model SPEAR yields a notable improvement on speech-related tasks over the audio-domain model SPEAR, as evidenced by tasks such as CD, SF-5, VI, and VL. We also observe a sharp increase for MST, a music transcription task, from SPEAR Large with 23.6 to SPEAR Large with 27.7. This suggests that the joint pre-training on speech and audio data enhances the model’s capability of performing fine-grained music tasks. Despite achieving a better overall score on HEAR, we do notice that the dual-domain model suffers from performance degradation in some environment and music-related tasks. This further motivates us to use a more balanced dataset containing more general audio data in future work.
@@
Appendix G Ablation Studies
@@
Teacher Model Selection: We compare pre-training with MVQ tokens extracted from different SSL teacher models (see Appendix G.1).
Masked Prediction Pre-training Loss: We investigate how to balance the pre-training loss on the masked and unmasked positions for SPEAR (see Appendix G.2).
Number of Codebooks: We study the effect of varying the number of codebooks in the MVQ quantiser on the pre-training performance (see Appendix G.3).
Feature Subspaces: We compare the feature subspaces reconstructed by the MVQ quantiser and a k-means clustering model (see Appendix G.4).
Dual-domain Pre-training: We investigate how the losses from speech and audio domains should be balanced in the dual-domain pre-training (see Appendix G.5).
G.1 Teacher model Selection
In this group of ablation studies, different choices of SSL models are compared for generating pre-training targets under the SPEAR framework. In addition to WavLM Large and Dasheng 1.2B as presented in Table 3, HuBERT-Large (Hsu et al., 2021) and ATST-frame (Li et al., 2024a), which are pre-trained with different SSL objectives and data scales, are used for generating fine-grained discrete targets for speech and audio data, respectively.
| Model | # Params | Targets |
SUPERB | ||||||||
| ASR | KS | IC | ASV | SID | SD | ER | ||||||
| Speech SSL Models | ||||||||||||
| HuBERT Large | 317M | - | 60k | 3.62 | 95.29 | 98.76 | 5.98 | 90.33 | 5.75 | 67.62 | ||
| WavLM Large | 317M | - | 94k | 3.44 | 97.86 | 99.31 | 3.77 | 95.49 | 3.24 | 70.62 | ||
| Ours | ||||||||||||
| Large-H-1 | 327M | HuBERT Large | 50k | 3.24 | 97.05 | 99.47 | 4.11 | 91.52 | 3.84 | 69.86 | ||
| Large-H-2 | 327M | HuBERT Large | 84k | 3.17 | 97.79 | 99.51 | 3.49 | 94.42 | 3.24 | 70.91 | ||
| SPEAR Large | 327M | WavLM Large | 84k | 3.27 | 97.89 | 99.47 | 3.14 | 95.49 | 3.20 | 71.88 | ||
@@ G.1 Teacher model Selection
HuBERT (Hsu et al., 2021) is a speech SSL model pre-trained using masked language modelling (MLM) loss on 60k hours of speech from Libri-light Kahn et al. (2020). In contrast to WavLM Large, HuBERT-Large is pre-trained on less diverse data (only read speech) without augmentations, resulting in weaker overall performance, especially on speaker-related tasks. A comprehensive comparison of HuBERT Large and WavLM Large on SUPERB can be found in Table 12 and Table 13.
Following Table 3, the representation from the 21st layer of HuBERT Large is used to train an MVQ quantiser with 16 codebooks for generating the pre-training targets. We train the following two Large models with pre-training targets generated from the HuBERT Large:
@@
Large-H-2: Pre-trained on Speech-84k using the same data augmentation as SPEAR Large, enabling a fair comparison with SPEAR Large and WavLM Large.
The pre-training performance is evaluated on SUPERB, and the results are shown in Table 16. As can be seen, MVQ tokens extracted from HuBERT Large are also effective pre-training targets. Large-H-1 outperforms its teacher HuBERT Large on all SUPERB tasks by a large margin on the premise of using the same amount of pre-training data. Notably, Large-H-1 demonstrates strong ASR performance, achieving the lowest WER on SUPERB ASR task, even surpassing SPEAR Large trained with more data, suggesting that the MVQ tokens extracted from HuBERT Large could have a stronger focus on semantic information (Mousavi et al., 2025).
By increasing the amount of pre-training data, Large-H-2 further improves over Large-H-1. However, Large-H-2 yields a weaker overall performance on SUPERB compared to SPEAR Large, which is pre-trained with MVQ tokens extracted from WavLM Large. This suggests that MVQ tokens extracted from stronger speech representations translate to a stronger per-training performance under SPEAR framework.
@@
ATST-frame (Li et al., 2024a) is an audio SSL model pre-trained with BYOL Grill et al. (2020) objective on 5k hours of AS-2M with 86M parameters. The model generates 768-d frame-level audio representations at a 25Hz frame rate. We train the following two Base models for comparison:
@@
| Model | # Params | Targets |
AudioSet | HEAR∗ | ||||||||
| AS-20k | AS-2M | ESC | FSD | GZ-Gen | NS-5 | LC | SC-5 | ||||||
| Audio SSL Models | |||||||||||||
| EAT (Chen et al., 2024) | 86M | - | 5k | 40.2 | 48.6 | - | - | - | - | - | - | ||
| ATST-frame | 88M | - | 5k | 39.0 | 48.0 | 89.0 | 55.7 | 88.3 | 68.6 | 78.1 | 92.6 | ||
| Ours | |||||||||||||
| Base-audio-1 | 94M | ATST-frame | 5k | 40.3 | 49.6 | 89.4 | 57.2 | 89.5 | 64.4 | 79.4 | 94.3 | ||
| Base-audio-2 | 94M | Dasheng 1.2B | 5k | 39.1 | 49.3 | 88.9 | 56.6 | 90.1 | 72.2 | 81.2 | 94.3 | ||
As can be seen in Table 17, Base-audio-1 exhibits very strong performance on AudioSet AT tasks, achieving higher mAP than its teacher ATST-frame. By yielding an mAP of 40.3 on AS-20k and 49.6 on AS-2M, Base-audio-1 outperforms EAT (Chen et al., 2024), setting a new state-of-the-art for audio SSL models pre-trained only on AS-2M. This validates the effectiveness of SPEAR as an audio SSL approach, as it shows that the student model can consistently outperform its teacher model used for generating the pre-training targets.
@@ Conclusion
G.2 Pre-training Loss
The hyperparameter in Equation 2 controls the contribution of the prediction loss on the masked and unmasked frames. To investigate its influence w.r.t SPEAR, we conducted experiments on single-domain models, varying from 0.0 (predicting only unmasked frames) to 1.0 (predicting only masked frames). The speech-domain models are pre-trained on LibriSpeech, while the audio models are pre-trained on AS-2M. The same MVQ tokens from Table 3 are used. We evaluate the downstream fine-tuning performance on LS-100 fine-tuning and AT, results presented in Figure 2.
@@ G.2 Pre-training Loss
G.3 Number of Codebooks
@@
Similar experiments are conducted for audio-domain pre-training. Following Table 3, the last layer of Dasheng 1.2B is used to train the MVQ quantiser with 4, 8, and 16 codebooks. The models are evaluated on the AudioSet fine-tuning task, and the results are shown in Table 19. As can be seen, increasing from 4 to 8 improves the downstream AT fine-tuning performance and the HEAR scores. However, further increasing to 16 degrades the pre-training performance, leading to a lower mAP and average HEAR score compared to . We hypothesise that representations of audio SSL models encapsulate less information compared to speech representations in general. Therefore, using a moderate number of codebooks seems to be enough for audio-domain pre-training. A too large might force some codebooks to capture the nuances in the audio teacher representations and introduce noise to the pre-training.
@@
G.4 Feature Subspaces of MVQ Tokens
In order to investigate if the codebooks in the MVQ quantiser have captured useful characteristics from the speech and audio representations, we visualise the reconstructed embedding space of the MVQ quantiser on a 2-D plane using UMAP (McInnes et al., 2018). Specifically, we visualise the speaker embeddings encoded by the MVQ quantiser of 10 speakers randomly drawn from LibriSpeech dev-clean sets, with each speaker having 25 utterances. The speech MVQ quantiser from Table 3 is used. The speaker embeddings are computed with the procedure described below. First, we use the speech MVQ quantiser (see Table 3) to encode the frame-level embeddings generated by WavLM Large into the MVQ tokens. Then, we compute the reconstructed frame-level embeddings by summing over the encoded code vector from each codebook. The speaker embedding for each utterance is obtained by calculating the mean embedding vector over all frames. As a comparison, we also visualize the speaker embedding space represented through a k-means clustering model. The 500-cluster k-means model used for generating the pre-training targets for WavLM Large is adopted, which is trained on the 9th layer representations of HuBERT Base. We use the cluster centroid to represent each frame and average the frame-level embeddings along the temporal dimension to obtain a single speaker embedding for each utterance.
The visualisations are shown in Figure 3(c). As can be seen, the MVQ quantiser successfully retains speaker characteristics, showing a clear separation between different speakers. It is noteworthy that a single codebook with 256 codes is capable of capturing certain levels of the speaker characteristics, showing reasonable separation between different speakers in Figure 3(b). However, the k-means centroids fail to distinguish different speakers, showing poor separation for different speakers. This observation aligns with the fact that our speech-domain model SPEAR Large achieves far better performance on ASV compared to WavLM Large, a task requiring distinguishing different speakers by comparing the speaker embedding similarity.
G.5 Dual-domain Pre-training
As mentioned in Section 3.2.2, we adopt an asymmetrical pre-training loss in Equation 6 for dual-domain pre-training. The following strategies for dual-domain pre-training are investigated:
Joint: Each training input induces two losses, computed against the and , regardless of the domain of .
Asymmetrical: For speech data, losses are computed against both and . For audio data, loss is only computed against . This approach is adopted by SPEAR.
We perform dual-domain pre-training using the Large size model on the Mix-97k data with the aforementioned three strategies. The models are evaluated after 100k training steps for quicker system verification, where we compare the results of LS-100 ASR fine-tuning, AS-20K AT fine-tuning, two SUPERB tasks (SID and PR), and the average score on HEAR. The results are shown in Table 20. Among the three strategies, Asymmetrical achieves a balanced performance across both domains. Computing the loss against the speech MVQ tokens for audio data is a useful regularisation to bridge the domain mismatch between speech and audio, preventing significant performance degradation on both domains (Asymmetrical vs Disjoint).
On the other hand, computing pre-training loss against the audio MVQ tokens for speech data in the Joint strategy is less useful. Compared to Asymmetrical strategy, an increase of 0.1 absolute WER and a 0.3 absolute lower average score on HEAR are observed. We suspect that this is caused by the imbalanced data distribution between speech and audio in Mix-97k, where speech data makes up 87% of the total data. The dominance of the speech data hinders the effective learning of generic audio representations for the Joint strategy. This motivates us to enlarge the proportion of general audio data in the total training corpora for our future work.
| Strategy | LS-100 | AS-20K | SUPERB | HEAR | ||
| test-clean | test-other | SID | PR | Avg | ||
| Joint | 2.9 | 5.9 | 36.8 | 90.6 | 3.24 | 78.7 |
| Disjoint | 3.0 | 5.8 | 37.0 | 87.4 | 3.40 | 78.3 |
| Asymmetrical | 2.9 | 5.8 | 36.9 | 90.7 | 3.12 | 79.0 |

### v2

@@ Appendix E HEAR Evaluation
| USAD Base | 94M | 95.8 | 80.0 | 93.6 | 82.2 | 52.2 | 94.0 | 100.0 | 86.3 | 78.7 | 26.7 | 97.3 | 95.7 | 81.6 | 57.0 | 96.6 | 97.6 | 19.5 | 76.0 | 78.4 |
| USAD Large | 330M | 94.1 | 79.5 | 93.9 | 83.4 | 53.0 | 87.4 | 100.0 | 97.6 | 79.1 | 38.4 | 97.4 | 96.1 | 83.2 | 57.0 | 97.0 | 97.5 | 18.5 | 75.3 | 79.4 |
| SPEAR, Base | 95M | 92.4 | 78.7 | 93.5 | 83.8 | 49.9 | 86.5 | 96.9 | 95.2 | 71.9 | 24.6 | 96.8 | 94.1 | 78.9 | 64.6 | 96.9 | 97.0 | 21.9 | 77.3 | 77.8 |
@@
| SPEAR, XLarge | 600M | 95.3 | 81.6 | 95.4 | 85.1 | 52.4 | 88.6 | 98.5 | 96.3 | 78.7 | 28.0 | 97.0 | 96.3 | 81.5 | 65.1 | 97.2 | 98.1 | 22.6 | 83.6 | 80.1 |
| SPEAR, Base | 95M | 95.3 | 82.1 | 94.9 | 85.9 | 54.2 | 88.8 | 100.0 | 95.2 | 76.9 | 39.7 | 97.1 | 96.0 | 82.6 | 69.4 | 97.3 | 98.2 | 24.6 | 85.6 | 80.6 |
| SPEAR, Large | 327M | 94.9 | 83.8 | 95.9 | 87.4 | 56.4 | 89.2 | 99.2 | 97.6 | 78.9 | 40.0 | 97.4 | 97.5 | 85.3 | 70.2 | 98.1 | 98.3 | 25.7 | 88.5 | 81.8 |
| SPEAR, XLarge | 600M | 95.8 | 83.6 | 96.0 | 89.4 | 57.1 | 91.0 | 100.0 | 96.9 | 80.9 | 41.1 | 97.4 | 97.9 | 86.0 | 74.2 | 98.4 | 98.6 | 26.6 | 90.4 | 82.6 |
@@
By performing unified speech and audio pre-training that incorporates more speech data, the dual-domain model SPEAR yields a notable improvement on speech-related tasks over the audio-domain model SPEAR, as evidenced by tasks such as CD, SF-5, VI, and VL. We also observe a sharp increase for MST, a music transcription task, from SPEAR Large with 23.6 to SPEAR Large with 26.9. This suggests that the joint pre-training on speech and audio data enhances the model’s capability of performing fine-grained music tasks. Despite achieving a better overall score on HEAR, we do notice that the dual-domain model suffers from performance degradation in some environments and music-related tasks. This further motivates us to use a more balanced dataset containing more general audio data in future work.
@@
Appendix F Ablation Studies
@@
Masked Prediction Pre-training Loss: We investigated how to balance the pre-training loss on the masked and unmasked positions for SPEAR (see Appendix F.2).
Encoder Architectures: We compared using Zipformer and Transformer as the encoder backbone for SPEAR (see Appendix F.3).
Number of Codebooks: We studied the effect of varying the number of codebooks in the MVQ quantiser on the pre-training performance (see Appendix F.5).
F.1 Teacher Models and Layers
F.1.1 Teacher Model Selection
In this group of ablation studies, different choices of SSL models are compared for generating pre-training targets under the SPEAR framework. In addition to WavLM Large and Dasheng 1.2B as presented in Section 4, HuBERT-Large (Hsu et al., 2021) and ATST-frame (Li et al., 2024), which are pre-trained with different SSL objectives and data scales, are used for generating fine-grained discrete targets for speech and audio data, respectively.
| Model | # Params | Targets | Pre-train Data | SUPERB | ||||||
| ASR | KS | IC | SV | SID | SD | ER | ||||
| Speech SSL Models | ||||||||||
| HuBERT Large | 317M | - | 60k | 3.62 | 95.29 | 98.76 | 5.98 | 90.33 | 5.75 | 67.62 |
| WavLM Large | 317M | - | 94k | 3.44 | 97.86 | 99.31 | 3.77 | 95.49 | 3.24 | 70.62 |
| Ours | ||||||||||
| Large-H-1 | 327M | HuBERT Large | 50k | 3.24 | 97.05 | 99.47 | 4.11 | 91.52 | 3.84 | 69.86 |
| Large-H-2 | 327M | HuBERT Large | 84k | 3.17 | 97.79 | 99.51 | 3.49 | 94.42 | 3.24 | 70.91 |
| SPEAR Large | 327M | WavLM Large | 84k | 3.27 | 97.89 | 99.47 | 3.14 | 95.55 | 3.2 | 72.1 |
@@ F.1.1 Teacher Model Selection
HuBERT (Hsu et al., 2021) is a speech SSL model pre-trained using masked language modelling (MLM) loss on 60k hours of speech from Libri-light (Kahn et al., 2020). In contrast to WavLM Large, HuBERT-Large is pre-trained on less diverse data (only read speech) without augmentations, resulting in weaker overall performance, especially on speaker-related tasks. A comprehensive comparison of HuBERT Large and WavLM Large on SUPERB can be found in Table 13 and Table 14. The representation from the 21st layer of HuBERT Large is used to train an MVQ quantiser with 16 codebooks for generating the pre-training targets. We train the following two Large models with pre-training targets generated from the HuBERT Large:
@@
Large-H-2: Pre-trained on Speech-84k using the WavLM-style data augmentation.
The pre-training performance is evaluated on SUPERB, and the results are shown in Table 17. Note that the SPEAR Large model in Table 17 is also pre-trained with WavLM-style data augmentation. As can be seen, MVQ tokens extracted from HuBERT Large are also effective pre-training targets. Large-H-1 outperforms its teacher HuBERT Large on all SUPERB tasks by a large margin on the premise of using the same amount of pre-training data. Notably, Large-H-1 demonstrates strong ASR performance, achieving the lowest WER on SUPERB ASR task, even surpassing SPEAR Large trained with more data, suggesting that the MVQ tokens extracted from HuBERT Large could have a stronger focus on semantic information (Mousavi et al., 2025).
By increasing the amount of pre-training data, Large-H-2 further improves over Large-H-1. However, Large-H-2 yields a weaker overall performance on SUPERB compared to SPEAR Large, which is pre-trained with MVQ tokens extracted from WavLM Large. This suggests that MVQ tokens extracted from stronger speech representations translate to a stronger pre-training performance under SPEAR framework.
@@
ATST-frame (Li et al., 2024) is an audio SSL model pre-trained with BYOL (Grill et al., 2020) objective on 5k hours of AS-2M with 86M parameters. The model generates 768-d frame-level audio representations at a 25Hz frame rate. We train the following two Base models for comparison:
@@
| Model | # Params | Targets | Pre-train Data | AudioSet | HEAR∗ | ||||||
| AS-20k | AS-2M | ESC | FSD | GZ-Gen | NS-5 | LC | SC-5 | ||||
| Audio SSL Models | |||||||||||
| EAT (Chen et al., 2024a) | 86M | - | 5k | 40.2 | 48.6 | - | - | - | - | - | - |
| ATST-frame | 88M | - | 5k | 39.0 | 48.0 | 89.0 | 55.7 | 88.3 | 68.6 | 78.1 | 92.6 |
| Ours | |||||||||||
| Base-audio-1 | 94M | ATST-frame | 5k | 40.3 | 49.6 | 89.4 | 57.2 | 89.5 | 64.4 | 79.4 | 94.3 |
| Base-audio-2 | 94M | Dasheng 1.2B | 5k | 39.1 | 49.3 | 88.9 | 56.6 | 90.1 | 72.2 | 81.2 | 94.3 |
As can be seen in Table 18, Base-audio-1 exhibits very strong performance on AudioSet AT tasks, achieving higher mAP than its teacher ATST-frame. By yielding an mAP of 40.3 on AS-20k and 49.6 on AS-2M, Base-audio-1 outperforms EAT (Chen et al., 2024a), setting a new state-of-the-art for audio SSL models pre-trained only on AS-2M. This validates the effectiveness of SPEAR as an audio SSL approach, as it shows that the student model can consistently outperform its teacher model used for generating the pre-training targets.
@@ Conclusion
F.1.2 Teacher Layer Selection
In this ablation study, experiments are carried out to compare different teacher layers for extracting pre-training targets. In our initial setup, the teacher layer is selected based on the downstream fine-tuning performance, i.e., ASR for speech teacher and AT for audio teacher. All experiments are conducted using the Base architecture, and the results are shown in Table 19.
For the speech teacher, we evaluate the 18th, 21st, and 24th (last) layers of WavLM Large and report the WERs on LS-100 fine-tuning tasks. The 21st layer is selected since it achieves the lowest WERs. Note that this choice aligns with the findings of Shi et al. (2024b), who also report the discrete tokens derived from the 21st layer of WavLM Large to yield strong ASR performance. For the audio teacher, we compare using the 25th, 35th, and 40th (last) layers of Dasheng-1.2B and report the mAP on the AS-20k fine-tuning task. The 40th layer is selected since it yields the highest mAP.
| WavLM layer | test-clean | test-other |
| 18 | 3.1 | 6.1 |
| 21 (adopted) | 3.0 | 5.8 |
| 24 | 3.1 | 6.0 |
| Dasheng layer | mAP |
| 25 | 38.4 |
| 35 | 38.7 |
| 40 (adopted) | 39.2 |
F.2 Pre-training Loss
The hyperparameter in Equation 1 controls the contribution of the prediction loss on the masked and unmasked frames. To investigate its influence w.r.t SPEAR, we conducted experiments on single-domain models, varying from 0.0 (predicting only unmasked frames) to 1.0 (predicting only masked frames). The speech-domain models are pre-trained on LibriSpeech, while the audio models are pre-trained on AS-2M. The same MVQ tokens from Section 4 are used. We evaluate the downstream fine-tuning performance on LS-100 fine-tuning and AT, results presented in Figure 3. Note that token mixing is not applied in this ablation.
@@ F.2 Pre-training Loss
F.3 Encoder Architecture
To decouple the influence of the encoder architecture from our pre-training framework, this ablation compares the Zipformer (utilising filterbank inputs) against a standard Transformer backbone (utilising waveform inputs). The Transformer implementation follows the architecture adopted in WavLM (Chen et al., 2022). We pre-train both models on the full LibriSpeech 960h corpus (LS-960) for 300k updates, maintaining the same MVQ-based pre-training objective. The results on LS-100 ASR fine-tuning and the SUPERB benchmark are presented in Table 20.
| Encoder Backbone | LibriSpeech finetune | SUPERB | |||||
| test-clean | test-other | PR | IC | KWS | SID | ER | |
| Transformer | 4.1 | 9.0 | 3.4 | 97.79 | 97.27 | 89.8 | 66.27 |
| Zipformer | 3.1 | 7.0 | 3.4 | 98.37 | 96.83 | 88.4 | 68.29 |
On the SUPERB tasks, which evaluate frozen representations, both architectures exhibit comparable performance: the Transformer proves stronger on SID, whilst the Zipformer excels on ER. However, the Zipformer demonstrates a distinct advantage in downstream ASR fine-tuning, a result consistent with its ASR-centric design (Yao et al., 2024). Moreover, the Zipformer is more computationally efficient due to its intermediate downsampling operations. For instance, pre-training the Zipformer requires approximately 350 GPU hours, roughly 60% of the 600 hours required for the Transformer. Consequently, the stronger ASR fine-tuning performance and computational efficiency motivate our selection of the Zipformer as the backbone architecture for SPEAR.
F.4 MVQ Tokens and k-means Tokens
F.4.1 Quantitative Evaluation
To isolate the impact of the quantization target, we conduct a controlled ablation comparing our MVQ tokens against standard k-means tokens. Both target types are derived from the same layer (the 21st layer) of WavLM Large, with the k-means baseline utilising 2000 clusters. Using the Base architecture, both models are pre-trained for 300k updates on LibriSpeech under identical augmentation strategies (noise mixing, utterance mixing, and masking). Notably, for the k-means experiment, we adopt the standard configuration by computing the loss solely on masked positions (Hsu et al., 2021). The results on the LS-100 ASR fine-tuning task and the SUPERB benchmark are shown in Table 21.
| Target | LibriSpeech finetune | SUPERB | |||||
| test-clean | test-other | PR | IC | KWS | SID | ER | |
| k-means, 2000 clusters | 3.5 | 7.2 | 4.0 | 97.92 | 96.79 | 86.6 | 67.56 |
| MVQ, 16 codebooks | 3.1 | 7.0 | 3.4 | 98.37 | 96.83 | 88.4 | 68.29 |
It can be seen that the model trained with MVQ tokens achieves better performance on all tasks, especially the two paralinguistic tasks: SID and ER. This aligns with our findings in Appendix F.4.2, where it is shown that the fine-grained MVQ tokens retain richer paralinguistic information than the coarse k-means tokens.
F.4.2 Feature Subspaces analysis
In order to investigate if the codebooks in the MVQ quantiser have captured useful characteristics from the speech and audio representations, we visualise the reconstructed embedding space of the MVQ quantiser on a 2-D plane using UMAP (McInnes et al., 2018). Specifically, we visualise the speaker embeddings encoded by the MVQ quantiser of 10 speakers randomly drawn from LibriSpeech dev-clean sets, with each speaker having 25 utterances. The speech MVQ quantiser from Sec 4 is used. The speaker embeddings are computed with the procedure described below. First, we use the speech MVQ quantiser (see Table 4) to encode the frame-level embeddings generated by WavLM Large into the MVQ tokens. Then, we compute the reconstructed frame-level embeddings by summing over the encoded code vector from each codebook. The speaker embedding for each utterance is obtained by calculating the mean embedding vector over all frames. As a comparison, we also visualize the speaker embedding space represented through a k-means clustering model. The 500-cluster k-means model used for generating the pre-training targets for WavLM Large is adopted, which is trained on the 9th layer representations of HuBERT Base. We use the cluster centroid to represent each frame and average the frame-level embeddings along the temporal dimension to obtain a single speaker embedding for each utterance.
The visualisations are shown in Figure 4. As can be seen, the MVQ quantiser successfully retains speaker characteristics, showing a clear separation between different speakers. It is noteworthy that a single codebook with 256 codes is capable of capturing certain levels of the speaker characteristics, showing reasonable separation between different speakers in Figure 4(b). However, the k-means centroids fail to distinguish different speakers, showing poor separation for different speakers. This observation aligns with the fact that our speech-domain model SPEAR Large achieves far better performance on SV compared to WavLM Large, a task requiring distinguishing different speakers by comparing the speaker embedding similarity.
F.5 Number of Codebooks
@@
Similar experiments are conducted for audio-domain pre-training. Following Section 4, the last layer of Dasheng 1.2B is used to train the MVQ quantiser with 4, 8, and 16 codebooks. The models are evaluated on the AudioSet fine-tuning task, and the results are shown in Table 23. As can be seen, increasing from 4 to 8 improves the downstream AT fine-tuning performance and the HEAR scores. However, further increasing to 16 degrades the pre-training performance, leading to a lower mAP and average HEAR score compared to . We hypothesise that representations of audio SSL models encapsulate less information compared to speech representations in general. Therefore, using a moderate number of codebooks seems to be enough for audio-domain pre-training. A too large might force some codebooks to capture the nuances in the audio teacher representations and introduce noise to the pre-training.
F.6 Loss Weighting
| LS-100 | AS-20k | ||
| test-clean | test-other | mAP | |
| 0.3 | 3.0 | 5.9 | 37.0 |
| 0.2 | 3.0 | 5.7 | 36.9 |
| 0.1 | 2.9 | 5.6 | 36.9 |
As shown in Section 3.2.2, the hyperparameter controls the contribution of the general-audio masked-prediction loss during joint training. To determine the optimal balance, we conduct an ablation study with 3 values of using our SPEAR Large architecture and compare the fine-tuning performance on LS-100 and AS-20k after 100k pre-training steps. The experimental results are shown in Table 24. We observed that reducing from 0.3 to 0.1 yields a 0.3 absolute WER improvement on test-other, while the mAP is only reduced by 0.1 absolute. Consequently, we adopted =0.1 in our dual-domain experiments for a balanced performance across both domains.
Appendix G Comparison with USAD
In this section, we performed a controlled comparison between SPEAR and USAD (Chang et al., 2025), another framework for joint speech and audio representation learning, also leveraging multiple domain-specific teachers. Specifically, we trained a new dual-domain SPEAR model with the Base architecture, named SPEAR (USAD-aligned), mirroring the USAD settings. We used the same teacher models as used in USAD, namely WavLM Base+ (speech) and ATST-Frame (Li et al., 2024) (audio) to extract the MVQ tokens as pre-training targets in SPEAR, aiming to rule out the effect of stronger teacher models being used in SPEAR. We also used a subset of the USAD training corpora, excluding Fisher and VoxLingua for speech and SoundNet for audio, due to availability issues. A summary of the model configurations and data usage for both models is shown in Table 25.
| Model | # Params | Speech data | Audio data | Total data |
| SPEAR (USAD-aligned) | 94M | 86k | 9.3k | 95.3k |
| USAD Base | 94M | 91k | 35k | 126k |
| Model | # Params | Data | SUPERB | HEAR | ||||||||
| PR | ASR | IC | KS | SID | ER | Env | Speech | Music | Avg | |||
| SPEAR (USAD-aligned) | 94M | 95.3k | 4.6 | 5.1 | 98.7 | 97.4 | 89.2 | 69.4 | 81.1 | 76.9 | 80.5 | 79.4 |
| USAD Base | 94M | 126k | 5.1 | 7.7 | 98.3 | 97.1 | 88.6 | 68.0 | 80.7 | 73.7 | 79.3 | 77.8 |
The comparison between SPEAR (USAD aligned) and USAD on SUPERB and HEAR is shown in Table 26. As can be seen, SPEAR (USAD-aligned) consistently outperforms USAD Base, despite only using a subset of the training data used by USAD, suggesting that SPEAR is more effective than USAD for learning unified speech and audio representations. We attribute this performance gap to the following two reasons:
Training objectives: In SPEAR, the student is trained to predict the discrete tokens extracted from teacher models given a masked input, which is a frequently used pretext task for SSL. This combination of KD and SSL in SPEAR enables the student to learn generic representations while benefiting from the knowledge of the two domain-specific teachers, creating a student model even with the capability of surpassing teacher models (e.g. SPEAR Large outperforms WavLM Large). On the other hand, USAD enforces the student to mimic the teacher representations through L1 and cosine distance loss. Consequently, the student performance is theoretically upper-bounded by the teacher’s performance.
Joint feature matching is ill-defined for disparate domains: In USAD, the student effectively minimises the distance to two embedding spaces (speech and audio) simultaneously. The L1 losses induced by two teachers encourage the student model to find a “mean” of two representation spaces. Since the feature spaces could be distinct, this “mean” representation may lie in a region of the manifold that lacks semantic meaning for either domain. SPEAR avoids this risk by quantising representations into discrete tokens via MVQ, where the tokens exhibit the capability of representing a certain characteristic of the input speech/audio data (see Appendix F.4.2, where we found a single codebook to contain rich speaker information). This allows the model to retain distinct, high-fidelity details by learning to predict the discrete tokens for both domains simultaneously, with lower risk of destructive interference.